Shift Summary

Nickolas Solomey (solomey@hep.uchicago.edu)
Tue, 13 Jul 1999 15:44:58 -0500

Shift Summary for day shift, 07/13/99

On Shift: Nick and Rick C.

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 45 min for laser run

Accounting of lost time: Beam down all day for maintance of Accelerator
========================================================================
Runs ended this shift:
13871 Long laser scan.
13872 Cosmic run, no tape writing, just for debuggin purposes.

Tapes brought to FCC this shift: QKx128-130 (where x=C,K,S), and QKB044
========================================================================
Accesses during shift:
7:00 Elizabeth and Ed make access to work on the HA problem.
7:15 Elizabeth and Ed come out of access with a bad HA power supply,
needs to be worked on today and put back in once repaired.
9:10 Dave + 2 other made access, checked the vacuum problem, first the
Blower B1 is completly dead, not fixable. Second,
the DP-1 overtemp protection switch had to be bypassed by jumpering
this switch out of circuit so there is now no over temp protection.
9:30 Angela, Colin W., and Marj make access to work on the DCs, Angela
took the gas sample and Marj and Colin worked on DC3y current
draw.
10:45 Ed drops interlock, but it was intentional.
11:50 Bud and Carrie remove Blower B1 from building for repairs, maybe
it can be fixed at factory in a week is Doug Jensen's hope.
12:30 Elizabeth work for 1 hour on changing the bad PMT base in BA.
14:30 Rick Ford and Coleman inspect NM2 for radiation hot spots,
none found, small dripping water leak on NM2Q1-3, reported.
14:35 Rick C looked at 90 deg Acc counter signals with no beam,
just target residual radiation, all counter signals look OK,
nim logic also looks OK. Connected 90 Acc 1,2,3 discriminator
signals to L1 spare hardlines so we can look at signals upstairs
when beam returns. May have to retreat from 1/2" hole to 1/4" hole.
========================================================================
Other activities during shift: Accelerator off all day for maintance day.
========================================================================
7:25 Run B-field of NM4AN down to low with the script AN832NEGLOW so
we are ready for a muon run when beam returns.
8:05 Took HA power supply to Bob Jones in high rise for repair, need
to pick up and install before beam returns since there is no spare.
11:00 Sydney has finished with upgrade for run control.
11:30 Marj increase gas flow to DC3 to see is inpurities or Helium
is causing DC3 current draw out of the spill.
15:00 Ed, Rick, Julie and Sydney look at CsI pin-diode problem in
the array, it comes and goes.
15:30 Picked up the HA power supply from being repaired.